OpenStudy (anonymous):

I need to get 400,000 (i wish) if i was to get 6% compounded annually what is the formula i would have to use? thanks in advance

OpenStudy (anonymous):

\[A=P\left(1+{r\over100}\right)^n\]

OpenStudy (anonymous):

amount needed = A = 400,000 principal = P = compute r = APR = 6 n = number of years for the think to happen

OpenStudy (anonymous):

hows that now?

OpenStudy (anonymous):

yes. could you chceck the question. some information is missing.

OpenStudy (anonymous):

already? by noon?

OpenStudy (anonymous):

yea. thats how long they make us wait! so, n = 80-55=25

OpenStudy (anonymous):

yep. and nobody tells us about the 401k. have to find out things the hard ways

OpenStudy (anonymous):

like the song goes, "my education came from reading the road-signs.."

OpenStudy (anonymous):

thats what you calculate. you've got A, r and n \[400000=P\left(1+{6\over100}\right)^{25}\]

OpenStudy (anonymous):

\[400000=P(1.06)^{25}\\ 400000=P(4.292)\]

OpenStudy (anonymous):

I dont see the need for a table!?

OpenStudy (anonymous):

but for what? you mean for the power(exponent)?

OpenStudy (anonymous):

I am in the dark here. I do not know what the tables represent and for what

OpenStudy (anonymous):

can you take a picture of the tables and attach here so I can take a look?

OpenStudy (anonymous):

P is a little less than $95K

OpenStudy (anonymous):

so, an investment of about 95K will give us 400K after 25 years

OpenStudy (anonymous):

but what is the number "1.4185"?

OpenStudy (anonymous):

so, all good?